# Trigonometry of partially truncated hyperbolic triangles and tetrahedra

### v2

@@ 4.1. One horoball
Using the parametrization (1) with $\mathbf{v}_{1}$ as starting point and tangent vector $-\mathbf{y}_{1}$ (since $\mathbf{y}_{1}$ is outward-pointing from $H_{1}$), we obtain $\tilde{\lambda}_{12}(s)=\cosh s\,\mathbf{v}_{1}-\sinh s\,\mathbf{y}_{1}$, an arclength parametrization for $\tilde{\lambda}_{12}$ having $\tilde{\lambda}_{12}(0)=\mathbf{v}_{1}$ and with $\tilde{\lambda}_{12}(s)\in H_{1}$ for small $s>0$. By Lemma 1.8, $\tilde{\lambda}_{34}$ is parametrized by $\tilde{\lambda}_{34}(t)=e^{-h_{34}}\cosh t\,\mathbf{x}_{4}+e^{-t}\mathbf{y}_{3}$.
@@
From this we see that “$D$”, now taken to represent a family of functions of $(s,t)$ parametrized by $(L_{12},e^{h_{34}},L_{13},L_{23},e^{h_{14}},e^{h_{24}})$, increases pointwise with each of the last four quantities for any fixed $(s,t)$; therefore its absolute minimum $T(x,y;L_{13},L_{23};e^{h_{14}},e^{h_{24}})$ does as well. ∎
@@ 4.2. Two horoballs
For $i=1,2$, let $\mathbf{y}_{i}$ be the outward normal to the half-space $H_{i}$ bounded by $P_{i}$ that contains $P_{3-i}$. Then $\mathbf{y}_{1}$ and $\mathbf{y}_{2}$ are oppositely-oriented tangent vectors to the geodesic intersecting $P_{1}$ and $P_{2}$ perpendicularly, so by Lemma 1.9 $\mathbf{y}_{1}\circ\mathbf{y}_{2}<0$ and $-\mathbf{y}_{1}\circ\mathbf{y}_{2}=L_{12}\doteq\cosh\ell_{12}$, where $\ell_{12}$ is the shortest distance between $P_{1}$ and $P_{2}$. By Lemma 1.8, the geodesic $\tilde{\lambda}_{13}$ perpendicular to $P_{1}$ and in the direction of $\mathbf{x}_{3}$ has the parametrization $\tilde{\lambda}_{13}(s)=e^{-h_{13}}\cosh s\mathbf{x}_{3}+e^{-s}\mathbf{y}_{1}$, where the signed distance $h_{13}$ from $P_{1}$ to $B_{3}$ satisfies $e^{h_{13}}=-\mathbf{y}_{1}\circ\mathbf{x}_{3}$ (noting that $\mathbf{y}_{1}\circ\mathbf{x}_{3}<0$, since $\mathbf{x}_{3}$ is in the ideal boundary of $H_{1}$ by hypothesis). The hypothesis and Lemma 1.8 likewise imply that for any $i\in\{1,2\}$ and $j\in\{3,4\}$, the signed distance $h_{ij}$ from $P_{i}$ to $B_{j}$ satisfies $e^{h_{ij}}=-\mathbf{y}_{i}\circ\mathbf{x}_{j}$, and that the geodesic $\tilde{\lambda}_{24}$ perpendicular to $P_{2}$ and in the direction of $\mathbf{x}_{4}$ has the parametrization $\tilde{\lambda}_{24}(t)=e^{-h_{24}}\cosh t\mathbf{x}_{4}+e^{-t}\mathbf{y}_{2}$. Lemma 1.4 implies that the signed distance $d_{34}$ from $B_{3}$ to $B_{4}$ satisfies $e^{d_{34}}=-\frac{1}{2}\mathbf{x}_{3}\circ\mathbf{x}_{4}$, so $-\mathbf{x}_{3}\circ\mathbf{x}_{4}=2e^{d_{34}}$.
@@
and $C_{--}=C_{+-}+C_{-+}-C_{++}+L_{12}$. It is clear that $0<C_{++}<C_{+-},C_{-+}<C_{--}$, so by Lemma 3.3, $D(s,t)$ has a unique critical point $(s_{0},t_{0})$ at which it attains an absolute minimum; and moreover, that each coordinate of this critical point is positive. This establishes the Proposition’s assertion (1), and (3) follows from the fact that each of $\tilde{\lambda}_{12}(s)$ and $\tilde{\lambda}_{34}(t)$ are parametrized pointing into $\Delta$ at 00.
@@
Using the parametrization (1) with $\mathbf{v}_{1}$ as starting point and tangent vector $-\mathbf{y}_{1}$ (since $\mathbf{y}_{1}$ is outward-pointing from $H_{1}$), we obtain $\tilde{\lambda}_{12}(s)=\cosh s\,\mathbf{v}_{1}-\sinh s\,\mathbf{y}_{1}$, an arclength parametrization for $\tilde{\lambda}_{12}$ having $\tilde{\lambda}_{12}(0)=\mathbf{v}_{1}$ and with $\tilde{\lambda}_{12}(s)\in H_{1}$ for small $s>0$.
@@
Let $C_{++}$ be the coefficient of $e^{s}e^{t}$ in the expression for $D(s,t)$ above, and name the coefficients of the other addends as $C_{+-}$, $C_{-+}$, $C_{--}$ correspondingly. We have:
@@ 4.4. Four horoballs
  • In the case that all $\mathbf{x}_{i}$ lie in a single plane, ie. if the ideal tetrahedron they determine degenerates to a quadrilateral, in which the line through $\mathbf{x}_{1}$ and $\mathbf{x}_{2}$ separates $\mathbf{x}_{3}$ from $\mathbf{x}_{4}$, applying Penner’s “ideal Ptolemy theorem” [9, Prop. 2.6(a)] to the formula gives $\cosh T_{4}(x,y;a,b,c,d)=1$, so the value of $T_{4}$ is 00. This is correct, reflecting that in this case the edges $\lambda_{12}$ and $\lambda_{34}$ are crossing diagonals.
  • The transversal length of a regular ideal tetrahedron with a maximal, fully symmetric horoball packing at its vertices is determined by the formula as $\cosh T_{4}(1,1;1,1,1,1)=2$. This can be confirmed by the angle of parallelism (see eg. [1, Th. 7.9.1(ii)]), since by symmetry considerations the transversal is the compact edge of a hyperbolic triangle with angles 00, $\pi/2$, and $\pi/6$.
Remarks.
@@
Note that $T_{4}$ is also invariant under transforming $(a,b,x)\mapsto t(a,b,x)$ for any $t\in(0,\infty)$. This records the effect on the inputs to $T_{4}$ of rescaling $\mathbf{x}_{1}$; ie. of making a different choice of horoball centered at the same ideal point $[\mathbf{x}_{1}]$. Correspondingly, $T_{4}$ is invariant under transformations $(c,d,x)\mapsto u(c,d,x)$, $(a,c,y)\mapsto v(a,c,y)$, and $(b,d,y)\mapsto w(b,d,y)$ for any $u$, $v$, or $w\in(0,\infty)$.

### v3

@@ 4.1. One horoball
Using the parametrization (1) with $\mathbf{v}_{1}$ as starting point and tangent vector $\mathbf{y}_{1}$ (since $\mathbf{y}_{1}$ is outward-pointing from $H_{1}$), we obtain $\tilde{\lambda}_{12}(s)=\cosh s\,\mathbf{v}_{1}-\sinh s\,\mathbf{y}_{1}$, an arclength parametrization for $\tilde{\lambda}_{12}$ having $\tilde{\lambda}_{12}(0)=\mathbf{v}_{1}$ and with $\tilde{\lambda}_{12}(s)\in H_{1}$ for small $s>0$. By Lemma 1.8, $\tilde{\lambda}_{34}$ is parametrized by $\tilde{\lambda}_{34}(t)=e^{-h_{34}}\cosh t\,\mathbf{x}_{4}+e^{-t}\mathbf{y}_{3}$.
@@
From this we see that “DD”, now taken to represent a family of functions of $(s,t)$ parametrized by $(L_{12},e^{h_{34}},L_{13},L_{23},e^{h_{14}},e^{h_{24}})$, increases pointwise with each of the last four quantities for any fixed $(s,t)$; therefore its absolute minimum $T(x,y;L_{13},L_{23};e^{h_{14}},e^{h_{24}})$ does as well. ∎
@@ 4.2. Two horoballs
For $i=1,2$, let $\mathbf{y}_{i}$ be the outward normal to the half-space $H_{i}$ bounded by $P_{i}$ that contains $P_{3-i}$. Then $\mathbf{y}_{1}$ and $\mathbf{y}_{2}$ are oppositely-oriented tangent vectors to the geodesic intersecting $P_{1}$ and $P_{2}$ perpendicularly, so by Lemma 1.9 $\mathbf{y}_{1}\circ\mathbf{y}_{2}<0$ and $\mathbf{y}_{1}\circ\mathbf{y}_{2}=L_{12}\doteq\cosh\ell_{12}$, where $\ell_{12}$ is the shortest distance between $P_{1}$ and $P_{2}$. By Lemma 1.8, the geodesic $\tilde{\lambda}_{13}$ perpendicular to $P_{1}$ and in the direction of $\mathbf{x}_{3}$ has the parametrization $\tilde{\lambda}_{13}(s)=e^{-h_{13}}\cosh s\mathbf{x}_{3}+e^{-s}\mathbf{y}_{1}$, where the signed distance $h_{13}$ from $P_{1}$ to $B_{3}$ satisfies $e^{h_{13}}=-\mathbf{y}_{1}\circ\mathbf{x}_{3}$ (noting that $\mathbf{y}_{1}\circ\mathbf{x}_{3}<0$, since $\mathbf{x}_{3}$ is in the ideal boundary of $H_{1}$ by hypothesis). The hypothesis and Lemma 1.8 likewise imply that for any $i\in\{1,2\}$ and $j\in\{3,4\}$, the signed distance $h_{ij}$ from $P_{i}$ to $B_{j}$ satisfies $e^{h_{ij}}=-\mathbf{y}_{i}\circ\mathbf{x}_{j}$, and that the geodesic $\tilde{\lambda}_{24}$ perpendicular to $P_{2}$ and in the direction of $\mathbf{x}_{4}$ has the parametrization $\tilde{\lambda}_{24}(t)=e^{-h_{24}}\cosh t\mathbf{x}_{4}+e^{-t}\mathbf{y}_{2}$. Lemma 1.4 implies that the signed distance $d_{34}$ from $B_{3}$ to $B_{4}$ satisfies $e^{d_{34}}=-\frac{1}{2}\mathbf{x}_{3}\circ\mathbf{x}_{4}$, so $\mathbf{x}_{3}\circ\mathbf{x}_{4}=2e^{d_{34}}$.
@@
and $C_{--}=C_{+-}+C_{-+}-C_{++}+L_{12}$. It is clear that $0<C_{++}<C_{+-},C_{-+}<C_{--}$, so by Lemma 3.3, $D(s,t)$ has a unique critical point $(s_{0},t_{0})$ at which it attains an absolute minimum; and moreover, that each coordinate of this critical point is positive. This establishes the Proposition’s assertion (1), and (3) follows from the fact that each of $\tilde{\lambda}_{12}(s)$ and $\tilde{\lambda}_{34}(t)$ are parametrized pointing into $\Delta$ at 0.
@@
Using the parametrization (1) with $\mathbf{v}_{1}$ as starting point and tangent vector $\mathbf{y}_{1}$ (since $\mathbf{y}_{1}$ is outward-pointing from $H_{1}$), we obtain $\tilde{\lambda}_{12}(s)=\cosh s\,\mathbf{v}_{1}-\sinh s\,\mathbf{y}_{1}$, an arclength parametrization for $\tilde{\lambda}_{12}$ having $\tilde{\lambda}_{12}(0)=\mathbf{v}_{1}$ and with $\tilde{\lambda}_{12}(s)\in H_{1}$ for small $s>0$.
@@
Let $C_{++}$ be the coefficient of $e^{s}e^{t}$ in the expression for $D(s,t)$ above, and name the coefficients of the other addends as $_{+-}, C+C_{-+}$, $C_{--}$ correspondingly. We have:
@@ 4.4. Four horoballs
  • In the case that all $\mathbf{x}_{i}$ lie in a single plane, ie. if the ideal tetrahedron they determine degenerates to a quadrilateral, in which the line through $\mathbf{x}_{1}$ and $\mathbf{x}_{2}$ separates $\mathbf{x}_{3}$ from $\mathbf{x}_{4}$, applying Penner’s “ideal Ptolemy theorem” [10, Prop. 2.6(a)] to the formula gives $\cosh T_{4}(x,y;a,b,c,d)=1$, so the value of $T_{4}$ is 0. This is correct, reflecting that in this case the edges $\lambda_{12}$ and $\lambda_{34}$ are crossing diagonals.
  • The transversal length of a regular ideal tetrahedron with a maximal, fully symmetric horoball packing at its vertices is determined by the formula as $\cosh T_{4}(1,1;1,1,1,1)=2$. This can be confirmed by the angle of parallelism (see eg. [1, Th. 7.9.1(ii)]), since by symmetry considerations the transversal is the compact edge of a hyperbolic triangle with angles 0, $\pi/2$, and $\pi/6$.
Remarks 1.
@@
Note that $T_{4}$ is also invariant under transforming $(a,b,x)\mapsto t(a,b,x)$ for any $t\in(0,\infty)$. This records the effect on the inputs to $T_{4}$ of rescaling $\mathbf{x}_{1}$; ie. of making a different choice of horoball centered at the same ideal point $\mathbf{x}_{1}]$. Correspondingly, $T_{4}$ is invariant under transformations $(c,d,x)\mapsto u(c,d,x)$, $(a,c,y)\mapsto v(a,c,y)$, and $(b,d,y)\mapsto w(b,d,y)$ for any $u$, $v$, or $w\in(0,\infty)$.